OpenStudy (anonymous):

Write equation of the line in slope intercept form that passes through the two points (-6,-3) (-8,4)

OpenStudy (luigi0210):

Easy, find the slope. then plug that into the point-slope formula and solve. \[ {\big \smile } \]

OpenStudy (anonymous):

again slope=-7/2 so equation of line is y=-7/2x-24

OpenStudy (luigi0210):

Ignore that "big smile" thing xD

OpenStudy (anonymous):

lol big smile

OpenStudy (luigi0210):

@quantun I hope you aren't doing the users work for them

OpenStudy (anonymous):

the slope is -7/2 so i take one of those points and plug them in for y and x in the formula

OpenStudy (luigi0210):

Correct. I'm sure you already know how he found the slope right?

OpenStudy (anonymous):

yes y2-y1/x2-x1

OpenStudy (luigi0210):

Good :)

OpenStudy (anonymous):

kk so it should look like 4=-7/2*8+b

OpenStudy (luigi0210):

Where are your x and y's?

OpenStudy (anonymous):

o shoot i wrote it wrong

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

4=-7/2 -8+b y=mx+b is that right?

OpenStudy (luigi0210):

First you use the point slope formula, then you change it to slope-intercept form

OpenStudy (luigi0210):

\[y-y1=m(x-x1)\]

OpenStudy (anonymous):

75.6

OpenStudy (anonymous):

wait i thought we already had the slope -7/2

OpenStudy (skullpatrol):

Let's start again please :-)

OpenStudy (luigi0210):

Yes, now just plug the slope and any point you choose into the equation.

OpenStudy (anonymous):

o lol

OpenStudy (anonymous):

so y- -3=-7/2(x- -6)

OpenStudy (luigi0210):

Yes, now just simplify :)

OpenStudy (anonymous):

kk

OpenStudy (skullpatrol):

y- (-3)=-7/2(x- (-6))

OpenStudy (anonymous):

i got y-(-3)=-7/2x+21 is that right and i just add the 3 to both side get rid of it right?

OpenStudy (luigi0210):

Hm, you made an error with your signs.

OpenStudy (anonymous):

o it should be a negative 21

OpenStudy (luigi0210):

Yup, and when you have a --, it'd be easier to just change them into a + right away so it doesn't confuse you

OpenStudy (anonymous):

o kk lol so make the negative 3 a positive 3 and subtract from both sides

OpenStudy (luigi0210):

Bingo!

OpenStudy (anonymous):

i got y=-7/2+8

OpenStudy (anonymous):

it will be -24

OpenStudy (luigi0210):

8? xD

OpenStudy (luigi0210):

Oh Miss Jessica :3

OpenStudy (anonymous):

lol i was typing to fast and hit 8 sry

OpenStudy (anonymous):

i meant -24 :)

OpenStudy (luigi0210):

Hm, well alright Good job!

OpenStudy (anonymous):

thank you everybody

OpenStudy (luigi0210):

You may have been here for a bit already but...

OpenStudy (luigi0210):

\[\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile } \]

OpenStudy (anonymous):

lol what is that

OpenStudy (anonymous):

O thx XD

OpenStudy (anonymous):

its really cool